OpenStudy (anonymous):

what is the chance it will rain 2 days out the week if the probability of rain is 0.2 percent

OpenStudy (kropot72):

The binomial distribution applies: \[P(2\ days\ out\ of\ the\ week)=\left(\begin{matrix}7 \\ 2\end{matrix}\right)0.2^{2}(1-0.2)^{5}=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (kropot72):

You're welcome :)